# Accelerating Image Generation with Sub-path Linear Approximation Model

### v2

@@ 5 Experiments
In this section, we conduct experiments to examine the performance of our proposed Sub-path Linear Approximation Model (SLAM). Firstly, we describe the experiment configuration and implementation details, and evaluate our models comprehensively on the text-to-image task (Sec. 5.1). Secondly, we verify the effectiveness of our algorithm design through detailed ablation studies (Sec. 5.2). Finally, we present the qualitative results of our SLAM. (Fig. 4).
@@ 5.1.1 Experimental Configuration
For evaluation, we adopt the commonly used FID and CLIP Score metrics. The results are reported on both SDv1.5 and SDv2.1-base backbones, thus verifying the generalizability of our method. For the experiment of distilling SDv2.1-base, we bench-mark our model on two test sets, LAION-Aesthetics-6+ as used in LCM [26] and MSCOCO2014-30k for zero-shot generalization. We also reproduce a SDv2.1-base LCM according to the training configuration outlined in [26] while replacing the $w$-condition with the fixed guidance scale, which has also improved the performance. We generally set the guidance scale for distilling SDv2.1-base to 8 and skipping step size to 20, which is consistent with  [26]. For the experiment of distilling SDv1.5, we compare our model with state-of-the-art generative models including foundation diffusion models, GANs, and accelerated diffusion models. The guidance scale is set to $3$ to obtain the optimal FID, and we adopt the huber [43] loss for our SLAD metric. The skipping step size is set to 100 for SLAM which has shown fast convergence. We examine our method on two commonly used benchmarks, MSCOCO2014-30k and MSCOCO2017-5k. More implementation details are provided in the supplementary materials.
@@ 5.1.2 Main Results
The results for SDv2.1-base are presented in Tab. 1, we use DDIM [42], DPM [23], DPM++ [24] and LCM [26] as baselines. It reveals that our SLAM surpasses baseline methods nearly across both test sets, at each step, and on both FID and CLIP Score metrics. We suppose that the close results on LAION are caused by overfitting, since the test set and train set are sourced from the same data collection. For SDv1.5 under the guidance scale $w=3$, the quantitative results are demonstrated in Tab. 2(a) and Tab. 2(c). Our model with 4 steps gets FID-30k of 10.06 and FID-5k of 20.77, which outperforms all other accelerated diffusion models, including flow-based method InstaFlow [21] and techniques that introduce GAN objectives such as UFOGen [49] and DMD [50]. Furthermore, SLAM showcases commensurate results with state-of-the-art foundation generative models such as DALL-E2 [32]. Even in two steps, SLAM has achieved a competitive performance of FID-30k 12.31 with parallel algorithms. In practical scenarios, a higher guidance scale $w$ is typically favored to enhance the resultant image quality. Accordingly, we trained our SLAM with $w$ set to 8 and bench-mark it against a range of advanced diffusion methodologies, as delineated in Tab. 2(c). In this regime, SLAM also demonstrates significant advantages, achieving state-of-the-art performance with a four-step FID-30k of 13.39 which exceeds other models by a large margin and is close to the teacher model. Notably, the FID-30k of our model with only two steps reaches 14.50, surpassing the four-step LCM and DMD. While DMD training consumes over one hundred A100 GPU days, which is more than 16 times our training duration.
@@ 5.2 Ablation Study
Figure 2: (a) Ablations on skipping step size and skipping mechanism. ME denotes for our Multiple Estimation strategy. (b) Training curve comparing LCM and SLAM. Our SLAM with step size 100 is conducted with ME, which brings faster convergence. (c) Estimation of the error $\delta$ between consistency mapping values of two adjacent points through PF-ODE. SLAM consistently outperforms LCM in terms of the error.
@@
Figure 3: (a) Visualization for different guidance scale $w$ on SLAM. (b) The trade-off curve of applying difference guidance scale. $w$ increases from $\{3.0,5.0,8.0,12.0\}$.
@@
Skipping Step Size & Training Cost Fig. 2(a) ablates the skipping step size during training, where we compare SLAM with or without the multiple estimation strategy (Sec. 4.2.2) and LCM. We can observe that: 1) Without multiple estimation, when the skipping step size $k$ is increasing, LCM suffers a more drastic decline in performance due to heightened optimization challenges for sub-path learning. Through leveraging our proposed Sub-path Linear ODE, SLAM can progressively learn the $dist_{\Delta}$ and effectively alleviate this collapse. 2) Equipped with the multiple estimation strategy, SLAM is capable of stably maintaining high image fidelity with large steps. Moreover, we compare the convergence trends between our method and LCM during training, as depicted in Fig. 2(b). When $k=20$, although our metrics initially converge more slowly during the early stages, the performance of our method gradually surpasses LCM by a large margin. It indicates that our training strategy provides a more effective learning objective, enabling SLAM to achieve a better result, while LCM quickly becomes overfitted. As $k$ raised to 100, larger skipping step size brings SLAM faster convergence that needs just 2K to 6K iterations which requires about only 6 A100 GPU days training, facilitating practical applications with fewer resources. Note that LCM needs 10k+ iterations for optimal performance which costs about 16 A100 GPU days and can not be applied to larger skipping step size due to the serious performance gap.
@@
Approximated Error Estimation for SLAM. To illustrate the efficacy of our approach, we directly estimate the denoising mapping error between two adjacent samples on the PF-ODE: $\delta(t,k)=\mathbb{E}[|\bm{f}_{\bm{\theta}}(\bm{x}_{t_{n+k},t_{n+k}}),\bm{f}_{\bm{\theta}}(\bm{x}_{t_{n}},t_{n}))|]$, which is firstly defined in Eq. 6. The results are shown in Fig. 2(c). We randomly selected 1000 samples from the COCO dataset and simulated adjacent points on the ODE by adding the same noise with adjacent timesteps. We utilize $k=20$ and the corresponding 50 timesteps for the DDIM scheduler, disregarding steps smaller than 100 due to their relatively larger simulation deviation. It can be seen that, especially at larger timesteps, the error $\delta$ of our SLAM is further reduced (about $10\%$ at $t=800$). This observation substantiates that SLAM indeed contributes to minimizing approximated errors, boosting the model’s capacity for high-quality image generation.
@@
The Effect of Guidance Scale $w$. The guidance scale $w$ is a critical hyper-parameter in Stable Diffusion [34, 10], with its adjustment allowing users to alter the semantic alignment and the quality of the generated image. In this study, we also examine the impact of varying the guidance scale $w$ for our SLAM based on SDv1.5, which is visualized in Fig. 3. As well as vanilla Stable Diffusion, while a higher $w$ value contributes to better sample quality as reflected by CLIP Scores, it concurrently leads to a degradation in FID performance and oversaturation.
@@
Figure 4: Comparsion of our SLAM and LCM [26] in 1,2 and 4-step generation. The results of LCM are based on our reproduction as illustrated in Sec. 5.1. SLAM has generated consistently higher-quality images that are clearer and more detailed. Noteworthy is the remarkable performance of SLAM in the 2-step generation, which aligns closely with the 4-step generation results of LCM, highlighting the efficiency and effectiveness of our approach in producing high-fidelity images with fewer generation steps.
@@
Figure 5: Qualitative Results. The text prompts are selected from DMD [50] in (a) and UFOGEN [49] in (b), and the results of the two are also cited from respective papers. Clearly, SLAM demonstrates the best generation quality in 4-step generation except for the SD models. When decreasing the sampling step to 2, SLAM still maintains a comparable performance, which generates even better results than 4-step LCM [26].
@@ 5.3 Qualitative Results
To emphasize the boosted generation quality of our SLAM, we display the 1,2 and 4-step generation results with the detailed comparison to LCM [26] in Fig. 4. Moreover, we compare our SLAM distilled from SDv1.5 [34] with the most advanced accelerating diffusion models in Fig. 5, which demonstrate that our SLAM has achieved the best generation quality across the existing methods.
@@ 6 Conclusion
In this paper, we propose a novel approach Sub-path Linear Approximation Models (SLAM) for accelerating diffusion models. SLAM leverages the approximation strategy in consistency models and considers the PF-ODE trajectories as a series of interconnected sub-paths delineated by sampled points. Guided by the optimization direction charted by each sub-path, Sub-path Linear (SL) ODEs also enable our approach to progressively and continuously optimize the approximated learning objectives and thus construct the denoising mappings with smaller cumulative errors. We also develop an efficient distillation procedure for SLAM to enable the incorporation of latent diffusion models. Extensive experiments on LAION, MS COCO 2014 and MS COCO 2017 datasets have consistently demonstrated the superiority of our method across existing accelerating diffusion approaches in a few-step generation with a fast training convergence.
@@ Appendix 0.A Implementation Details
For SDv2.1-base: To make a fair comparison between LCM and our approach, we train our SLAM with skipping step size $k=20$ and L2 distance metric function, which is consistent with [26] (the multiple estimation strategy is disabled since $k\leq k_{\phi}$). The difference is that we trained our models with a fixed guidance scale $w=8$. We also reproduce an SDv2.1-base LCM according to the training configuration outlined in [26] while replacing the $w$-condition with the fixed guidance scale, which has also improved the performance. For training costs, one GPU day can process 1.6K iterations with batch size of 1024 and without multiple estimation (ME) strategy. We train SLAM on SDv2.1-base with 80k iterations, which costs about 60 A100 GPU days.
For SDv1.5: The guidance scale $w$ is set to $3$ to obtain the optimal FIDs, and we adopt the huber loss in [43] as we find that it provides faster convergence and better performance on two-step and one-step. Guidance scale $w=8$ is also conducted and all ablation studies are under SDv1.5 with $w=8$ by default. For training costs, we train SLAM on SDv1.5 with multiple estimation (ME) described in Sec. 4.2 in the main paper, which executes solver $\Phi$ for $\frac{k}{k_{\phi}}$ times in one iteration. In our platform, when $k$=100 and $k_{\phi}=20$, one GPU day can process nearly half the amount of data compared to previously. As the skipping step size $k$ set to $100$ has shown fast convergence, it just costs 6k training iterations which require only 6 GPU days.
@@ Appendix 0.D Additional Generated Results
Since LCM has only released one versionhttps://huggingface.co/SimianLuo/LCM_Dreamshaper_v7 of SDv1.5 based model that is distilled from DreamShaper-v7, we train a SLAM on DreamShaper-v7 as well and compare it to LCM, as shown in Fig. 6 and Fig. 7.
Refer to captionRefer to captionRefer to caption
Figure 6: Generated samples on 2-step from our SLAM and LCM distilled from DreamShaper-v7. We leverage the LCM-Dreamshaper-v7 checkpoint hosted on Hugging Face.
Refer to captionRefer to captionRefer to caption
Figure 7: Generated samples on 4-step from our SLAM and LCM distilled from DreamShaper-v7. We leverage the LCM-Dreamshaper-v7 checkpoint hosted on Hugging Face.
Figure 8: Generated samples from SLAM (our method) and LCM.
Figure 9: Generated samples from SLAM (our method) and LCM.
Figure 10: Generated samples from SLAM (our method) and LCM.
@@
Here we provide more generated images from our SLAM, which are compared with LCM [26], the newly released one-step generation approach InstaFlow [21] and Stable Diffusion [34]. Unless otherwise specified, our SLAM’s generated results are based on the model distilled from the open source SDv1.5, and the results for LCM are generated from our reproduction. The images are shown in Figs. 8, 9, 10, 11 and 12.
Figure 11: Comparison of our method SLAM with one-step generation (InstaFlow [21]) and Stable Diffusion [34].
Figure 12: Comparison of our method SLAM with one-step generation (InstaFlow [21]) and Stable Diffusion [34].

### v3

@@ 5 Experiments
In this section, we conduct experiments to examine the performance of our proposed Sub-Path Linear Approximation Model (SPLAM). Firstly, we describe the experiment configuration and implementation details, and evaluate our models comprehensively on the text-to-image task (Sec. 5.1). Secondly, we verify the effectiveness of our algorithm design through detailed ablation studies (Sec. 5.2). Finally, we present the qualitative results of our SPLAM. (Sec. 5.3).
@@ 5.1.1 Experimental Configuration
For evaluation, we adopt the commonly used FID and CLIP Score metrics. The results are reported on both SDv1.5 and SDv2.1-base backbones, thus verifying the generalizability of our method. For the experiment of distilling SDv2.1-base, we bench-mark our model on two test sets, LAION-Aesthetics-6+ as used in LCM [26] and MSCOCO2014-30k for zero-shot generalization. We also reproduce a SDv2.1-base LCM according to the training configuration outlined in [26] while replacing the $w$-condition with the fixed guidance scale, which has also improved the performance. We generally set the guidance scale for distilling SDv2.1-base to 8 and skipping step size to 20, which is consistent with  [26]. For the experiment of distilling SDv1.5, we compare our model with state-of-the-art generative models including foundation diffusion models, GANs, and accelerated diffusion models. The guidance scale is set to $3$ to obtain the optimal FID, and we adopt the huber [43] loss for our SPLAD metric. The skipping step size is set to 100 for SPLAM which has shown fast convergence. We examine our method on two commonly used benchmarks, MSCOCO2014-30k and MSCOCO2017-5k. More implementation details are provided in the supplementary materials.
@@ 5.1.2 Main Results
The results for SDv2.1-base are presented in Tab. 1, we use DDIM [42], DPM [23], DPM++ [24] and LCM [26] as baselines. It reveals that our SPLAM surpasses baseline methods nearly across both test sets, at each step, and on both FID and CLIP Score metrics. We suppose that the close results on LAION are caused by overfitting, since the test set and train set are sourced from the same data collection. For SDv1.5 under the guidance scale $w=3$, the quantitative results are demonstrated in Tab. 2(a) and Tab. 2(c). Our model with 4 steps gets FID-30k of 10.06 and FID-5k of 20.77, which outperforms all other accelerated diffusion models, including flow-based method InstaFlow [21] and techniques that introduce GAN objectives such as UFOGen [49] and DMD [50]. Furthermore, SPLAM showcases commensurate results with state-of-the-art foundation generative models such as DALL-E2 [32]. Even in two steps, SPLAM has achieved a competitive performance of FID-30k 12.31 with parallel algorithms. In practical scenarios, a higher guidance scale $w$ is typically favored to enhance the resultant image quality. Accordingly, we trained our SPLAM with $w$ set to 8 and bench-mark it against a range of advanced diffusion methodologies, as delineated in Tab. 2(c). In this regime, SPLAM also demonstrates significant advantages, achieving state-of-the-art performance with a four-step FID-30k of 13.39 which exceeds other models by a large margin and is close to the teacher model. Notably, the FID-30k of our model with only two steps reaches 14.50, surpassing the four-step LCM and DMD. While DMD training consumes over one hundred A100 GPU days, which is more than 16 times our training duration.
@@ 5.2 Ablation Study
Figure 2: (a) Ablations on skipping step size and skipping mechanism. ME denotes for our Multiple Estimation strategy. (b) Training curve comparing LCM and SPLAM. Our SPLAM with step size 100 is conducted with ME, which brings faster convergence. (c) Estimation of the error $\delta$ between consistency mapping values of two adjacent points through PF-ODE. SPLAM consistently outperforms LCM in terms of the error.
@@
Figure 3: (a) Visualization for different guidance scale $w$ on SPLAM. (b) The trade-off curve of applying difference guidance scale. $w$ increases from $\{3.0,5.0,8.0,12.0\}$.
@@
Skipping Step Size & Training Cost Fig. 2(a) ablates the skipping step size during training, where we compare SPLAM with or without the multiple estimation strategy (Sec. 4.2.2) and LCM. We can observe that: 1) Without multiple estimation, when the skipping step size $k$ is increasing, LCM suffers a more drastic decline in performance due to heightened optimization challenges for sub-path learning. Through leveraging our proposed Sub-Path Linear ODE, SPLAM can progressively learn the $dist_{\Delta}$ and effectively alleviate this collapse. 2) Equipped with the multiple estimation strategy, SPLAM is capable of stably maintaining high image fidelity with large steps. Moreover, we compare the convergence trends between our method and LCM during training, as depicted in Fig. 2(b). When $k=20$, although our metrics initially converge more slowly during the early stages, the performance of our method gradually surpasses LCM by a large margin. It indicates that our training strategy provides a more effective learning objective, enabling SPLAM to achieve a better result, while LCM quickly becomes overfitted. As $k$ raised to 100, larger skipping step size brings SPLAM faster convergence that needs just 2K to 6K iterations which requires about only 6 A100 GPU days training, facilitating practical applications with fewer resources. Note that LCM needs 10k+ iterations for optimal performance which costs about 16 A100 GPU days and can not be applied to larger skipping step size due to the serious performance gap.
@@
Approximated Error Estimation for SPLAM. To illustrate the efficacy of our approach, we directly estimate the denoising mapping error between two adjacent samples on the PF-ODE: $\delta(t,k)=\mathbb{E}[|\bm{f}_{\bm{\theta}}(\bm{x}_{t_{n+k},t_{n+k}}),\bm{f}_{\bm{\theta}}(\bm{x}_{t_{n}},t_{n}))|]$, which is firstly defined in Eq. 6. The results are shown in Fig. 2(c). We randomly selected 1000 samples from the COCO dataset and simulated adjacent points on the ODE by adding the same noise with adjacent timesteps. We utilize $k=20$ and the corresponding 50 timesteps for the DDIM scheduler, disregarding steps smaller than 100 due to their relatively larger simulation deviation. It can be seen that, especially at larger timesteps, the error $\delta$ of our SPLAM is further reduced (about $10\%$ at $t=800$). This observation substantiates that SPLAM indeed contributes to minimizing approximated errors, boosting the model’s capacity for high-quality image generation.
@@
The Effect of Guidance Scale $w$. The guidance scale $w$ is a critical hyper-parameter in Stable Diffusion [34, 10], with its adjustment allowing users to alter the semantic alignment and the quality of the generated image. In this study, we also examine the impact of varying the guidance scale $w$ for our SPLAM based on SDv1.5, which is visualized in Fig. 3. As well as vanilla Stable Diffusion, while a higher $w$ value contributes to better sample quality as reflected by CLIP Scores, it concurrently leads to a degradation in FID performance and oversaturation.
@@
Figure 4: Comparsion of our SPLAM and LCM [26] in 1,2 and 4-step generation. The results of LCM are based on our reproduction as illustrated in Sec. 5.1. SPLAM has generated consistently higher-quality images that are clearer and more detailed. Noteworthy is the remarkable performance of SPLAM in the 2-step generation, which aligns closely with the 4-step generation results of LCM, highlighting the efficiency and effectiveness of our approach in producing high-fidelity images with fewer generation steps.
@@
Figure 5: Qualitative Results. The text prompts are selected from DMD [50] in (a) and UFOGEN [49] in (b), and the results of the two are also cited from respective papers. Clearly, SPLAM demonstrates the best generation quality in 4-step generation except for the SD models. When decreasing the sampling step to 2, SPLAM still maintains a comparable performance, which generates even better results than 4-step LCM [26].
@@ 5.3 Qualitative Results
To emphasize the boosted generation quality of our SPLAM, we display the 1,2 and 4-step generation results with the comparison to LCM [26] in Fig. 4. Moreover, we compare our SPLAM distilled from SDv1.5 [34] with the most advanced accelerating diffusion models in Fig. 5, which demonstrate that our SPLAM has achieved the best generation quality across the existing methods.
@@ 6 Conclusion
In this paper, we propose a novel approach Sub-Path Linear Approximation Models (SPLAM) for accelerating diffusion models. SPLAM leverages the approximation strategy in consistency models and considers the PF-ODE trajectories as a series of interconnected sub-paths delineated by sampled points. Guided by the optimization direction charted by each sub-path, Sub-Path Linear (SL) ODEs also enable our approach to progressively and continuously optimize the approximated learning objectives and thus construct the denoising mappings with smaller cumulative errors. We also develop an efficient distillation procedure for SPLAM to enable the incorporation of latent diffusion models. Extensive experiments on LAION, MS COCO 2014 and MS COCO 2017 datasets have consistently demonstrated the superiority of our method across existing accelerating diffusion approaches in a few-step generation with a fast training convergence.
Acknowledgments
This work is supported by the National Key R&\&&D Program of China (No. 2022ZD0160900), the National Natural Science Foundation of China (No. 62076119, No. 61921006), the Fundamental Research Funds for the Central Universities (No. 020214380119), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.
@@ Appendix 0.A Implementation Details
For SDv2.1-base: To make a fair comparison between LCM and our approach, we train our SPLAM with skipping step size $k=20$ and L2 distance metric function, which is consistent with [26] (the multiple estimation strategy is disabled since $k\leq k_{\phi}$). The difference is that we trained our models with a fixed guidance scale $w=8$. We also reproduce an SDv2.1-base LCM according to the training configuration outlined in [26] while replacing the $w$-condition with the fixed guidance scale, which has also improved the performance. For training costs, one GPU day can process 1.6K iterations with batch size of 1024 and without multiple estimation (ME) strategy. We train SPLAM on SDv2.1-base with 80k iterations, which costs about 60 A100 GPU days.
For SDv1.5: The guidance scale $w$ is set to $3$ to obtain the optimal FIDs, and we adopt the huber loss in [43] as we find that it provides faster convergence and better performance on two-step and one-step. Guidance scale $w=8$ is also conducted and all ablation studies are under SDv1.5 with $w=8$ by default. For training costs, we train SPLAM on SDv1.5 with multiple estimation (ME) described in Sec. 4.2 in the main paper, which executes solver $\Phi$ for $\frac{k}{k_{\phi}}$ times in one iteration. In our platform, when $k$=100 and $k_{\phi}=20$, one GPU day can process nearly half the amount of data compared to previously. As the skipping step size $k$ set to $100$ has shown fast convergence, it just costs 6k training iterations which require only 6 GPU days.
@@ Appendix 0.D Additional Generated Results
Since LCM has only released one versionhttps://huggingface.co/SimianLuo/LCM_Dreamshaper_v7 of SDv1.5 based model that is distilled from DreamShaper-v7, we train a SPLAM on DreamShaper-v7 as well and compare it to LCM, as shown in Fig. 6 and Fig. 7.
Figure 6: Generated samples on 2-step from our SPLAM and LCM distilled from DreamShaper-v7. We leverage the LCM-Dreamshaper-v7 checkpoint hosted on Hugging Face.
Figure 7: Generated samples on 4-step from our SPLAM and LCM distilled from DreamShaper-v7. We leverage the LCM-Dreamshaper-v7 checkpoint hosted on Hugging Face.
Figure 8: Generated samples from SPLAM (our method) and LCM.
Figure 9: Generated samples from SPLAM (our method) and LCM.
@@
Here we provide more generated images from our SPLAM, which are compared with LCM [26], the newly released one-step generation approach InstaFlow [21] and Stable Diffusion [34]. Unless otherwise specified, our SPLAM’s generated results are based on the model distilled from the open source SDv1.5, and the results for LCM are generated from our reproduction. The images are shown in Figs. 8, 9, 10 and 11.
Figure 10: Comparison of our method SPLAM with one-step generation (InstaFlow [21]) and Stable Diffusion [34].
Figure 11: Comparison of our method SPLAM with one-step generation (InstaFlow [21]) and Stable Diffusion [34].